What is the mass (grams) of 4.62x1024 in molecules of glucose (C6H12O6)


1
Expert's answer
2021-04-30T07:42:27-0400

Solution:

One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

n(C6H12O6) = (4.62×1024 molecules of C6H12O6) × (1 mol / 6.022×1023 molecules) = 7.672 mol C6H12O6

The molar mass of C6H12O6 is 180.156 g/mol.

Hence,

(7.672 mol C6H12O6) × (180.156 g C6H12O6 / 1 mol C6H12O6) = 1382.15 g C6H12O6

Mass of C6H12O6 = 1382.15 g


Answer: 1382.15 grams of glucose (C6H12O6) are in 4.62×1024 molecules of glucose.